### v1

@@ 3.3 Exposure-Aware Fusion (EAF) Block
Our novel exposure aware fusion block is architecturally designed to integrate both local and global enhancement features, enabling comprehensive and cohesive image enhancement. The fusion process begins with two convolutional layers that apply spatial filtering to extract the salient features necessary for exposure correction. We also use global average pooling, mapping the feature maps to the global context vector. These fusion weights serve as a gating mechanism to regulate the contribution of local and global features. They are adaptively learned, encapsulating both detailed texture information and broad illumination context.
Table 1: Results for our exposure guided transformer approach over ME-v2 [2] and SICE-v2 [4] datasets. , , denotes top three respectively. We did not include other recent models that are too complex ($>10M$ params).
@@
Zero-DCE [15] 14.55 0.5887 10.40 0.5142 12.06 0.5441 16.92 0.6330 7.11 0.4292 12.02 0.5211 0.079M
Zero-DCE++ [34] 13.82 0.5887 9.74 0.5142 11.37 0.5583 11.93 0.4755 6.88 0.4088 9.41 0.4422 0.010M
DPED [24] 20.06 0.6826 13.14 0.5812 15.91 0.6219 16.83 0.6133 7.99 0.4300 12.41 0.5217 0.390M
KIND [73] 15.51 0.7115 11.66 0.7300 13.20 0.7200 15.03 0.6700 12.67 0.6700 13.85 0.6700 0.590M
DeepUPE [55] 19.10 0.7321 14.69 0.7011 16.25 0.7158 16.21 0.6807 11.98 0.5967 14.10 0.6387 7.790M
SID [7] 19.37 0.8103 18.83 0.8055 19.04 0.8074 19.51 0.6635 16.79 0.6444 18.15 0.6540 7.760M
SID-ENC [17] 22.59 0.8423 22.36 0.8519 22.45 0.8481 21.36 0.6652 19.38 0.6843 20.37 0.6748 >7.760M
RUAS [37] 13.43 0.6807 6.39 0.4655 9.20 0.5515 16.63 0.5589 4.54 0.3196 10.59 0.4394 0.002M
SCI [44] 9.96 0.6681 5.83 0.5190 7.49 0.5786 17.86 0.6401 4.45 0.3629 12.49 0.5051 0.001M
MSEC [2] 20.52 0.8129 19.79 0.8156 20.08 0.8210 19.62 0.6512 17.59 0.6560 18.58 0.6536 7.040M
CMEC [48] 22.23 0.8140 22.75 0.8336 22.54 0.8257 17.68 0.6592 18.17 0.6811 17.93 0.6702 5.400M
LCDPNet [54] 22.35 0.8650 22.17 0.8476 22.30 0.8552 17.45 0.5622 17.04 0.6463 17.25 0.6043 0.960M
DRBN [66] 19.74 0.8290 19.37 0.8321 19.52 0.8309 17.96 0.6767 17.33 0.6828 17.65 0.6798 0.530M
DRBN+ERL [20] 19.91 0.8305 19.60 0.8384 19.73 0.8355 18.09 0.6735 17.93 0.6866 18.01 0.6796 0.530M
DRBN-ERL+ENC [20] 22.61 0.8578 22.45 0.8724 22.53 0.8651 22.06 0.7053 19.50 0.7205 20.78 0.7129 0.580M
ELCNet [22] 22.37 0.8566 22.70 0.8673 22.57 0.8619 22.05 0.6893 19.25 0.6872 20.65 0.6861 0.018M
IAT [11] 20.34 0.8440 21.47 0.8518 20.91 0.8479 21.41 0.6601 22.29 0.6813 21.85 0.6707 0.090M
ELCNet+ERL [20] 22.48 0.8424 22.58 0.8667 22.53 0.8545 22.14 0.6908 19.47 0.6982 20.81 0.6945 0.018M
FECNet [19] 22.19 0.8562 23.22 0.8748 22.70 0.8655 22.01 0.6737 19.91 0.6961 20.96 0.6849 0.150M
FECNet+ERL [20] 23.10 0.8639 23.18 0.8759 23.15 0.8711 22.35 0.6671 20.10 0.6891 21.22 0.6781 >0.150M
@@
[…]
3.4 Loss Functions
To enhance multiple aspects of image quality, our training uses a detailed loss function setup in RGB color space. It includes $L_{1}$ and $L_{2}$ losses for handling outliers and detail, SSIM for structural integrity, and VGG for semantic consistency. We also incorporate a novel MUL-ADD (MA) loss to adjust the image’s contrast and brightness accurately, ensuring that the dynamic range is well represented without blurring details. The VGG loss helps match the output to high-level visual quality standards. Our combined loss function $\mathcal{L}_{total}$, considering both local and global outputs, is detailed below:
$\mathcal{C}_{total}(y,\hat{y},I_{low},I_{high})=\alpha L_{1}(y,\hat{y})_{(l,g)}+\beta L_{2}(y,\hat{y})+\gamma L_{SSIM}(y,\hat{y})$ (3)
$+~{}\delta L_{VGG}(y,\hat{y})+\eta L_{MA}(M_{L},A_{L},I_{low},I_{high}+L_{attn}(M_{g}(b),M)$ (4)
where $\alpha,\beta,\gamma,\delta,$ and $\eta$ are hyperparameters balancing the influence of each loss term, $y$ is the ground truth, $\hat{y}$ is the predicted image, $M_{L}$ and $A_{L}$ are the multiplicative and additive components of the local block, $I_{low}$ is the low-light input, and $I_{high}$ is the target high-quality image. Our fine-tuning stage’s loss equation can be presented with the physics based KL-divergence loss:
$\mathcal{L}_{finetune}(y,\hat{y},P,Q)=$ $\ \lambda L_{1}(y,\hat{y})+\mu L_{SSIM}(y,\hat{y})+\nu L_{KL}(\mathcal{P},Q)_{(l,g)}$ (5)
@@ 3.4 Loss Functions
Methods LOL-v1 LOL-v2 MIT-Fivek Average
DRBN [66] 19.55/0.746 20.13/0.820 -/- 19.84/0.783
DPE -/- -/- 23.80/0.880 -
Deep-UPE [55] -/- 13.27/0.452 23.04/0.893 18.15/0.672
3D-LUT [71] 16.35/0.585 17.59/0.721 25.21/0.922 19.71/0.742
DRBN+ERL [20] 19.84/0.824 -/- 22.14/0.873 20.99/0.848
ECLNet+ERL [20] 22.01/0.827 -/- 23.71/0.853 22.39/0.840
FECNet+ERL [20] 21.08/0.829 -/- 24.18/0.864 22.63/0.846
RetinexNet [60] 16.77/0.462 18.37/0.723 -/- 17.57/0.592
KinD++ [72] 21.30/0.823 19.08/0.817 -/- 20.19/0.820
EnlightenGAN [27] 17.483/0.652 18.64/0.677 -/- 18.06/0.664
IAT [11] 23.38/0.809 23.50/0.824 25.32/0.920 24.07/0.851
LLFormer [57] 25.75/0.823 26.19/0.819 -/- 25.97/0.816
MIRNet [70] 24.10/0.832 20.35/0.782 -/- 22.22/0.807
U-EGformer 23.56/0.836 22.05/0.841 24.89/0.928 23.5/0.869
(a) Results on LOL-v1 [60], LOL-v2 [66],
Adobe FiveK [3], and average.
@@
RetinexNet 12.397 0.606 0.407 12.450 0.619 0.364
ZeroDCE 12.428 0.633 0.362 12.475 0.644 0.314
RAUS 0.864 0.493 0.525 0.8628 0.494 0.499
SGZ 10.866 0.607 0.415 10.987 0.621 0.364
LLFlow 12.737 0.617 0.388 12.737 0.617 0.388
URetinexNet 10.903 0.600 0.402 10.894 0.610 0.356
SCI 8.644 0.529 0.511 8.559 0.532 0.484
KinD 12.986 0.656 0.346 13.144 0.668 0.302
KinD++ 13.196 0.657 0.334 13.235 0.666 0.295
@@
U-EGformer 14.724 0.665 0.269 15.101 0.670 0.260
(b) Results on SICE Grad and SICE Mix datasets
[…]
@@ 4.1 Framework Setting
Datasets.  In our study, we employ eight diverse datasets to rigorously train and evaluate our proposed model: LOL-v1 and LOL-v2 for foundational training and testing with real-world and synthetic images; Multiple-Exposure ME-v2 tailored for diverse exposure scenarios; SICE, including the SICE-Grad and SICE-Mix subsets for gradient and mixed-exposure challenges, respectively; and MIT-FiveK for benchmarking against professionally retouched images. LOL-v1 [60] contains 500 image pairs with 485 and 15 for training and testing datasets, where each image has a resolution of ($3\times 600\times 400$). LOL-v2 is divided into real and synthetic subsets with detailed configurations for training/testing (with 689 and 100 images for real-world); in the BAcklit Image Dataset (BAID) dataset [43], we only use 380 randomly selected training images from Liang et al. [36] and utilize the complete 368 2K resolution images from the test set.
Training Strategy.  In tackling the mixed exposure challenge in image processing, our approach adopts a pre-training stage and a finetuning stage as shown in Fig. 2. We engage in pre-training using our custom loss function, $\mathcal{C}_{total}$ (Eq. 3.4), which combines several loss components with individually set hyperparameters for input-output pairs. In the finetuning phase, we refine the model with a physics-based pixel-wise reconstruction loss function tailored to camera sensors obeying Poisson distribution $\mathcal{P}$ [59] (more details are in the supplementary material).
@@ 4.2 Qualitative Results
[…]
Figure 7: We show the SSIM-map comparison between IAT [11] and the proposed model.
[…]
Components  1 2 3 4 5 6 7 8
@@
LEB Attention map
GEB Attention map
(Inverse illuminance
map) * (input)
(Illuminance map)
* (input)
Attention Trans. Block 3 8 8 5 5 5 5
@@
PSNR 21.12 18.28 18.82 20.32 20.66 21.33 21.90 22.05
SSIM 81.56 72.12 74.24 81.96 81.36 79.86 83.92 84.10
[…]
The LOL dataset is still one of the challenging datasets even for state-of-the-art models due to its extremely low-light scenario. In Fig. 9. top, we compare recent top models, where most of the models fail to match the color of the wood pane in this case with a lower PSNR score. In Fig. 9. bottom, we show visual results for the SICE Grad dataset for the mixed-exposure task. In Fig. 13, we show a few most challenging examples where LEB and GEB alone could not manage certain cases with extreme low and bright pixels, where the EAF block helps in re-highlighting the attended features from GAMG.
[…]
@@
Ablation Study.  Our framework utilizes a data-centric approach with a smaller memory footprint ($\sim$12.5 Mb777over LOL-v2 input image) and computation alongside other strategies as we have shown through Tab. 1’s ‘#params’ column. Through Tab. 8, we show the effectiveness of each module in our framework over LOL-v2 dataset. We demonstrate that the inverse illuminance map, combined with the attention map and exposure-aware fusion block, achieves the best results when configured with the appropriate combination of loss functions. The first column achieves better performance on LOLv2. Additionally, Tab. 8 highlights a notable improvement in LPIPS score compared to the closest baseline [72].

### v2

@@ 3.3 Exposure-Aware Fusion (EAF) Block
Our novel exposure aware fusion block is architecturally designed to integrate local and global enhancement features, enabling comprehensive image enhancement. The fusion process begins with two convolutional layers that apply spatial filtering to extract the salient features necessary for exposure correction. We also use global average pooling, mapping the feature maps to the global context vector. These fusion weights serve as a gating mechanism to regulate the contribution of local and global features. They are adaptively learned, encapsulating both detailed texture information and broad illumination context.
3.4 Loss Functions
To enhance multiple aspects of image quality, our training uses a detailed loss function setup in the RGB color space. It includes $L_{1}$ and $L_{2}$ losses for handling outliers and preserving fine details, SSIM for maintaining structural integrity, and VGG for ensuring semantic consistency. We also incorporate a novel MUL-ADD loss to accurately adjust the image’s contrast and brightness, ensuring that the dynamic range is well represented without blurring details. The VGG loss helps match the output to high-level visual quality standards.
MUL-ADD loss. The Multiplicative-Additive loss is specifically designed to handle mixed exposure scenarios by optimizing the multiplicative and additive adjustments applied during the enhancement process. This loss is mathematically defined as:
$L_{\text{MA}}(M_{L},A_{L},I_{\text{low}},I_{\text{high}})=\boldsymbol{\xi}~L_{1}(M_{L},\hat{M}(I_{\text{low}},I_{\text{high}}))+\boldsymbol{\psi}~L_{1}(A_{L},\hat{A}(I_{\text{low}},I_{\text{high}}))\vskip-20.00003pt$ (3)
Where:
$\vskip-20.00003pt\hat{M}(I_{\text{low}},I_{\text{high}})=\frac{I_{\text{high}}-\hat{A}(I_{\text{low}},I_{\text{high}})}{I_{\text{low}}+\epsilon},\quad\hat{A}(I_{\text{low}},I_{\text{high}})=I_{\text{high}}-\frac{I_{\text{high}}\cdot\epsilon}{I_{\text{low}}+\epsilon}$ (4)
In these equations, $I_{\text{low}}$ and $I_{\text{high}}$ represent the low-light input and the target high-quality image, respectively, while $M_{\text{L}}$ and $A_{\text{L}}$ are the multiplicative and additive components learned by the local enhancement block. The parameters $\boldsymbol{\xi}$ and $\boldsymbol{\psi}$ control the balance between the two components of the Mul-Add loss. The small constant $\epsilon=1e-8$ is added to avoid division by zero. This loss function ensures that the multiplicative and additive factors are optimized to produce a natural-looking enhancement without overcompensating in either direction. Our combined loss function $\mathcal{\text{L}}_{\text{C}}$, considering both local and global outputs, is detailed below:
$\begin{split}\mathcal{L}_{\text{C}}(y,\hat{y},I_{low},I_{high})=&\ \boldsymbol{\alpha}~L_{1}(y,\hat{y})_{(l,g)}+\boldsymbol{\beta}~L_{2}(y,\hat{y})+\boldsymbol{\gamma}~L_{\text{SSIM}}(y,\hat{y})+\boldsymbol{\delta}~L_{\text{VGG}}(y,\hat{y})\\ &+\boldsymbol{\eta}~L_{\text{MA}}(M_{L},A_{L},I_{\text{low}},I_{\text{high}})+L_{\text{attn}}(M_{g}(b),M)\end{split}$ (5)
where $\alpha,\beta,\gamma,\delta,$ and $\eta$ are hyperparameters balancing the influence of each loss term, $y$ is the ground truth, $\hat{y}$ is the predicted image, $M_{L}$ and $A_{L}$ are the multiplicative and additive components of the local block, $I_{low}$ is the low-light input, and $I_{high}$ is the target high-quality image. Our fine-tuning stage’s loss equation can be presented with the physics-based KL-divergence loss:
$\mathcal{L}_{\text{finetune}}(y,\hat{y},P,Q)=$ $\ \boldsymbol{\lambda}~L_{1}(y,\hat{y})+\boldsymbol{\mu}~L_{\text{SSIM}}(y,\hat{y})+\boldsymbol{\nu}~L_{\text{KL}}(\mathcal{P},Q)_{(l,g)}$ (6)
Table 1: Results for our exposure guided transformer approach over ME-v2 [afifi2021learning] and SICE-v2 [Cai2018deep] datasets. , , denotes top three respectively. We did not include other recent models that are too complex (>10M>10M params).
@@ 3.4 Loss Functions
Zero-DCE [guo2020zero] 14.55 0.5887 10.40 0.5142 12.06 0.5441 16.92 0.6330 7.11 0.4292 12.02 0.5211 0.079M
Zero-DCE++ [Zero-DCE++] 13.82 0.5887 9.74 0.5142 11.37 0.5583 11.93 0.4755 6.88 0.4088 9.41 0.4422 0.010M
DPED [ignatov2017dslr] 20.06 0.6826 13.14 0.5812 15.91 0.6219 16.83 0.6133 7.99 0.4300 12.41 0.5217 0.390M
KIND [10.1145/3343031.3350926] 15.51 0.7115 11.66 0.7300 13.20 0.7200 15.03 0.6700 12.67 0.6700 13.85 0.6700 0.590M
DeepUPE [Wang_2019_CVPR] 19.10 0.7321 14.69 0.7011 16.25 0.7158 16.21 0.6807 11.98 0.5967 14.10 0.6387 7.790M
SID [ChenCXK18] 19.37 0.8103 18.83 0.8055 19.04 0.8074 19.51 0.6635 16.79 0.6444 18.15 0.6540 7.760M
SID-ENC [huang2022exposure] 22.59 0.8423 22.36 0.8519 22.45 0.8481 21.36 0.6652 19.38 0.6843 20.37 0.6748 >7.760M
RUAS [ll_benchmark] 13.43 0.6807 6.39 0.4655 9.20 0.5515 16.63 0.5589 4.54 0.3196 10.59 0.4394 0.002M
SCI [Ma_2022_CVPR] 9.96 0.6681 5.83 0.5190 7.49 0.5786 17.86 0.6401 4.45 0.3629 12.49 0.5051 0.001M
MSEC [afifi2021learning] 20.52 0.8129 19.79 0.8156 20.08 0.8210 19.62 0.6512 17.59 0.6560 18.58 0.6536 7.040M
CMEC [nsampi2021learning] 22.23 0.8140 22.75 0.8336 22.54 0.8257 17.68 0.6592 18.17 0.6811 17.93 0.6702 5.400M
LCDPNet [wang2022local] 22.35 0.8650 22.17 0.8476 22.30 0.8552 17.45 0.5622 17.04 0.6463 17.25 0.6043 0.960M
DRBN [drbn] 19.74 0.8290 19.37 0.8321 19.52 0.8309 17.96 0.6767 17.33 0.6828 17.65 0.6798 0.530M
DRBN+ERL [huang2023learning] 19.91 0.8305 19.60 0.8384 19.73 0.8355 18.09 0.6735 17.93 0.6866 18.01 0.6796 0.530M
DRBN-ERL+ENC [huang2023learning] 22.61 0.8578 22.45 0.8724 22.53 0.8651 22.06 0.7053 19.50 0.7205 20.78 0.7129 0.580M
ELCNet [huang2017arbitrary] 22.37 0.8566 22.70 0.8673 22.57 0.8619 22.05 0.6893 19.25 0.6872 20.65 0.6861 0.018M
IAT [Cui_2022_BMVC] 20.34 0.8440 21.47 0.8518 20.91 0.8479 21.41 0.6601 22.29 0.6813 21.85 0.6707 0.090M
ELCNet+ERL [huang2023learning] 22.48 0.8424 22.58 0.8667 22.53 0.8545 22.14 0.6908 19.47 0.6982 20.81 0.6945 0.018M
FECNet [10.1145/3343031.3350855] 22.19 0.8562 23.22 0.8748 22.70 0.8655 22.01 0.6737 19.91 0.6961 20.96 0.6849 0.150M
FECNet+ERL [huang2023learning] 23.10 0.8639 23.18 0.8759 23.15 0.8711 22.35 0.6671 20.10 0.6891 21.22 0.6781 >0.150M
@@
$\text{U-EGformer}_{\text{eaf}}$ 22.82 0.8578 22.90 0.8558 22.86 0.8568 22.98 0.7192 21.84 0.7102 22.41 0.7179 0.102M
@@
(a) Results on LOL-v1 [WeiWY018], LOL-v2 [drbn], Adobe FiveK [fivek]
Methods LOL-v1 LOL-v2 MIT-FiveK
PSNR $\uparrow$ SSIM $\uparrow$ PSNR $\uparrow$ SSIM $\uparrow$ PSNR $\uparrow$ SSIM $\uparrow$
DRBN [drbn] 19.55 0.746 20.13 0.820 - -
DPE - - - - 23.80 0.880
Deep-UPE [Wang_2019_CVPR] - - 13.27 0.452 23.04 0.893
3D-LUT [zeng2020lut] 16.35 0.585 17.59 0.721 25.21 0.922
DRBN+ERL [huang2023learning] 19.84 0.824 - - 22.14 0.873
ECLNet+ERL [huang2023learning] 22.01 0.827 - - 23.71 0.853
FECNet+ERL [huang2023learning] 21.08 0.829 - - 24.18 0.864
RetinexNet [WeiWY018] 16.77 0.462 18.37 0.723 - -
KinD++ [kind_plus] 21.30 0.823 19.08 0.817 - -
ElightenGAN [jiang2021enlightengan] 17.483 0.652 18.64 0.677 - -
IAT [Cui_2022_BMVC] 23.38 0.809 23.50 0.824 25.32 0.920
LLFormer [wang2023ultra] 25.75 0.823 26.19 0.819 - -
MIRNet [Zamir2020MIRNet] 24.10 0.832 20.35 0.782 - -
U-EGformer 23.56 0.836 22.05 0.841 24.89 0.928
(b) Results on SICE Grad and SICE Mix datasets [zheng2022low].
@@
RetinexNet [WeiWY018] 12.397 0.606 0.407 12.450 0.619 0.364
ZeroDCE [guo2020zero] 12.428 0.633 0.362 12.475 0.644 0.314
RAUS [liu2021ruas] 0.864 0.493 0.525 0.8628 0.494 0.499
SGZ [zheng2022semantic] 10.866 0.607 0.415 10.987 0.621 0.364
LLFlow [wang2021low] 12.737 0.617 0.388 12.737 0.617 0.388
URetinexNet [9879970] 10.903 0.600 0.402 10.894 0.610 0.356
SCI [ma2022toward] 8.644 0.529 0.511 8.559 0.532 0.484
KinD [10.1145/3343031.3350926] 12.986 0.656 0.346 13.144 0.668 0.302
KinD++ [kind_plus] 13.196 0.657 0.334 13.235 0.666 0.295
@@
U-EGformer (finetuned) 14.724 0.665 0.269 15.101 0.670 0.260
@@ 4.1 Framework Setting
Datasets. We employ eight diverse datasets to train and evaluate our proposed model: LOL-v1 [WeiWY018] and LOL-v2 for foundational training and testing with real-world and synthetic images; Multiple-Exposure ME-v2 tailored for diverse exposure scenarios; SICE, including the SICE-Grad and SICE-Mix subsets for gradient and mixed-exposure challenges, respectively; and MIT-FiveK for benchmarking against professionally retouched images. LOL-v1 contains 500 image pairs with 485 and 15 for training and testing datasets, where each image has a resolution of ($3\times 600\times 400$). LOL-v2 is divided into real and synthetic subsets with detailed configurations for training/testing (with 689 and 100 images for real-world); in the BAcklit Image Dataset (BAID) dataset [lv2022backlitnet], we only use 380 randomly selected training images from Liang et al. [liang2023iterative] and utilize the complete 368 2K resolution images from the test set.
Training Strategy.  In tackling the mixed exposure challenge in image processing, our approach adopts a pre-training stage and a finetuning stage as shown in Fig. 2. We engage in pre-training using our custom loss function, $\mathcal{L}_{C}$ (Eq. 5), which combines several loss components with individually set hyperparameters for input-output pairs. In the finetuning phase, we refine the model with a physics-based pixel-wise reconstruction loss function tailored to camera sensors obeying Poisson distribution $\mathcal{P}$ [wang2023exposurediffusion] (more details are in the supplementary material in Section 7.4).
@@ 4.2 Qualitative Results
Figure 8: A comparison of the SSIM-maps for IAT [Cui_2022_BMVC] and the proposed model.
Table 3: Visual results of our module settings in the pipeline over LOLv2-real dataset. “”(resp.“”) means the module was unused (resp. used). ‘*’ represents multiple warmup restarts. : $L_{1}+\text{MSE}+KL_{\mathcal{P}}+\text{SSIM}$: $L_{1}+\text{MSE}+\text{KL}_{\mathcal{P}}+\text{SSIM}+\text{VGG}$: $L_{1}+\text{MSE}+\text{KL}_{\mathcal{P}}+\text{SSIM}+\text{Mul-Add}$: $L_{1}+\text{MSE}+\text{KL}_{\mathcal{P}}+\text{SSIM}+\text{VGG}+\text{MA-SL}_{1}$
Components Settings
1 2 3 4 5 6 7 8
@@
LEB (Attention map)
GEB (Attention map)
GEB (Inverse Illuminance map + input)
GEB (Illuminance map + input)
Attention Transformer Block 3 8 8 5 5 5 5
@@
PSNR $\uparrow$ 21.12 18.28 18.82 20.32 20.66 21.33 21.90 22.05
SSIM $\uparrow$ 81.56 72.12 74.24 81.96 81.36 79.86 83.92 84.10
The LOL dataset is a challenging dataset even for state-of-the-art models due to its extremely low-light scenario. In Fig. 7 top, we compare recent top models, where most of the models fail to match the color of the wood pane in this case with a lower PSNR score. In Fig. 7 bottom, we show visual results for the SICE Grad dataset for the mixed-exposure task. In Supplementary Material (Fig. 12)we show a few challenging examples where LEB and GEB alone could not manage certain cases with extreme low and bright pixels, where the EAF block helps in re-highlighting the attended features.
Quantitative Comparison.  Tab. 1 reports PSNR and SSIM scores for U-EGformer and $\text{U-EGformer}_{eaf}$. U-EGformer demonstrates superior performance in handling both underexposure and overexposure scenarios across ME-v2 and SICE-v2 datasets, outperforming the majority of existing methods with significantly fewer parameters. Despite SID-L [huang2022exposure] utilizing >10M>10M parameters we perform quite similarly (with differences of 0.0398/0.05 for SSIM/PSNR), while U-EGformer is 115 times smaller network than SID-L. In Tab. 2(a), we show remarkable generalization across LOL-v1, LOL-v2, and MIT-FiveK datasets, outperforming many baselines and illustrating robustness in exposure correction. Moreover, Tab. 2(b) sets new benchmarks on the challenging SICE Grad and SICE Mix datasets, underscoring U-EGformer’s superior performance in correcting mixed exposure images. In Fig. 8, we show a direct comparison of SSIM maps over the enhanced outputs of IAT and U-EGformer. Darker pixels in SSIM maps as seen more in IAT than in U-EGformer, indicate areas where the enhanced outputs from the two frameworks significantly differ with ground-truth. Moreover, in Fig. 10, we demonstrate the noise and the color consistency that visually seems better in U-EGformer’s output.
@@
Ablation Study.  Our framework utilizes a data-centric approach with a smaller memory footprint ($\sim$12.5 Mb555over LOL-v2 input image) and computation alongside other strategies as we have shown through Tab. 1’s ‘#params’ column. Through Tab. 3, we show the effectiveness of each module in our framework over LOL-v2 dataset. We demonstrate that the inverse illuminance map, combined with the attention map and exposure-aware fusion block, achieves the best results when configured with the appropriate combination of loss functions. The first column achieves better performance on LOLv2.